# Bidirectional predictive coding

### v1

@@ 5 Conclusion
We begin with a detailed explanation of the methods underlying the experiments presented in the main paper. Next, we provide summary tables containing the exact values plotted in the figures. We then explore the learning performance of three additional predictive coding models. Following that, we examine the effect of scaling the discriminative and generative energies in bPC on both supervised and unsupervised learning performance. We also analyse the impact of removing max-pooling on generative artefacts. Finally, we present samples generated by a combined discBP and genBP model.
@@ Appendix A Details of experiments
Here, we detail the training setup and the evaluation procedure of the experiments in the paper. The code with all models and experiments will be available upon publication at https://github.com/Bogacz-Group. Our implementation of the predictive coding models was adapted from pcx_.
@@ A.1 Datasets
We evaluate the models on four standard image classification benchmarks: MNIST [lecun2010mnist_], Fashion-MNIST [xiao2017fashionmnist_], CIFAR-10, and CIFAR-100 [krizhevsky2009learning_]. Below, we summarise their properties and how they were subdivided in our experiments for training, validation and testing. The images of all datasets were normalised so that pixel values varied between $-1$ and $1$.
@@ A.3 Training procedures
Require : Model with neural activities $x$, parameters $\theta$, energy $E$ and initialisation function $init(\cdot)$. Dataset $\{y_{p}\}_{p=1}^{P}$ with $P$ mini-batches of $B$ elements. Number of epochs $N$. Activity optimiser $optim_{x}(\cdot)$, number of activity updates $K$, and parameter optimiser $optim_{\theta}(\cdot)$
for $n=1$ to $N$ do
@@
Require : Model with parameters $\theta$ and forward pass $fp(\cdot,\theta)$. Energy function $E$. Dataset $\{y_{p}\}_{p=1}^{P}$ with $P$ mini-batches of $B$ elements. Number of epochs $N$. Parameter optimiser $optim_{\theta}(\cdot)$
for $n=1$ to $N$ do
@@
The optimisation procedures differ between activity updates and parameter updates. In Algorithm 1, neural activities are updated using stochastic gradient descent with momentum. This optimiser outperforms Adam in PC models [pcx_]. In contrast, Algorithm 2 does not include activity updates and only optimises parameters. For parameter updates in both algorithms, we use the AdamW optimiser [loshchilov2017decoupled_]. The optimiser hyperparameters were determined via a hyperparameter search.
@@
The hybridBP model includes a stop-gradient operation in its energy function. This ensures that its second loss term is only applied during parameter updates, not iterative inference. This loss term objective is to improve the discriminative initialisation of the model’s latent layer. In the broader machine learning literature, hybridBP can be viewed as a generative model that performs Bayesian inference by combining amortised inference with iterative refinement of the latent variables [marino2018iterative_].
@@ A.4 Training hyperparameters
In this Section, we report the hyperparameter search space for the experiments in Sections 4.1, 4.2, 4.3, and 4.5, as well as the training parameters for the models in Section 4.4. All hyperparameter tuning was performed using Bayesian optimisation from Weights and Biases [wandb_].
Tables 5, 6, and 7 list the hyperparameter search spaces for the models trained in Sections 4.1, 4.2, and 4.3, respectively. The hyperparameter search was conducted using the validation sets of each dataset. In Section 4.1, discPC and genPC were tuned for classification accuracy and generation RMSE separately, while bPC and hybridPC were jointly tuned on both metrics. This ensures a fair comparison, as bPC and hybridPC have twice as many parameters as discPC and genPC. For joint tuning, we combined the two metrics using the objective $2\cdot(1-\text{accuracy}/100)+\text{RMSE}$. In Section 4.2, the models were tuned for the image reconstruction MSE, after which the RMSE was reported for consistency with the generation RMSE. In Section 4.3, all models were jointly optimised for classification accuracy and reconstruction MSE. We combined the two metrics using $2\cdot(1-\text{accuracy}/100)+\text{MSE}$ for MNIST and Fashion-MNIST, and $(1-\text{accuracy}/100)+4\cdot\text{MSE}$ for CIFAR-10. The RMSE was also reported after tuning instead of the MSE. The weightings in the combined objectives compensate for the scale difference between the metrics.
@@ A.5 Evaluation
Classification accuracy
We evaluate classification performance using accuracy, defined as the percentage of input images correctly classified by each model.
@@
Conditional generation of mean class images
We evaluate the generative performance of the models by computing the root mean squared error (RMSE) between images generated for each class and the corresponding class-average image. We obtain the class-average image by computing the average image across all images belonging to that class in the evaluation set. To compute the RMSE, we calculate the squared error for each pixel between the generated image and the average image for the corresponding class. These errors are averaged across all pixels and classes, and the square root of this mean is reported as the final RMSE.
@@
Image reconstruction from representations
We assess the quality of learned representations by measuring the RMSE between reconstructed images and their original inputs. RMSE is computed in the same manner as for conditional image generation.
@@
Visualisation of Energy Landscapes on XOR
@@
We visualise the energy landscapes of trained genPC, discPC, and bPC models on the XOR task. To generate the landscape, we clamp the input layer neurons to a 2D coordinate within the range [-3,3], sampled at intervals of 0.25 along both axes. Simultaneously, we clamp the output layer $x_{L}$ to the one-hot encoding of one of the two classes. For each coordinate–label combination, we run 10,000 steps of iterative inference to ensure convergence to equilibrium and record the final energy of the model. This procedure is repeated over the grid of 2D inputs and both class labels, allowing us to plot the full energy landscape. For fair comparison, we also include the combined energy landscape of discPC and genPC. Together, the combined parameter count of discPC and genPC matches the total parameter count of bPC. Equal weighting is applied to the energy values of the discPC and genPC models when adding their energy, as we use equal weighting for the discriminative and generative components in the bPC model ($\alpha_{disc}=\alpha_{gen}=1$).
MNIST image generation with PC models
@@
We evaluate the generative capabilities of the bPC model and a combined genPC and discPC model on the MNIST dataset. The goal is to assess how likely these models are to assign low energy to implausible, label-inconsistent samples.
@@
  • Fréchet Inception Distance (FID): The FID measures the distance between the distributions of real and generated images [HeuselRUNKH17_]. We use the public implementation of FID from Seitzer2020FID_, but modify it to use a ResNet classifier trained on MNIST, instead of the original Inception model trained on ImageNet. This ensures that the FID score better reflects visual quality and diversity on the MNIST domain.
  • Inception Score (IS): The Inception Score evaluates how easily a classifier can identify the class of a generated image [salimans2016improvedtechniquestraininggans_]. We used a publicly available implementation for the MNIST dataset [inception_].
@@
To assess model robustness to missing input data, we evaluate classification accuracy under varying levels of input occlusion. This evaluation follows the same procedure as standard classification (described above), but for images missing a random subset of pixels. The proportion of missing pixels ranges from 10% to 90%, in increments of 10%. Missing pixels are selected uniformly at random, independently of their spatial location within the image. For all models, missing pixels in $x_{1}$ are initialised to zero activity and are left unclamped, allowing the model to update their values during inference. We repeat the same classification procedure as before, but we increased the number of inference steps to 600,000 to ensure convergence. This is necessary because predictive coding models converge more slowly when input information is incomplete [tang2023recurrent_]. In addition to measuring classification accuracy, we also record the post-inference activity of neurons in $x_{1}$ to qualitatively assess the model’s ability to fill in the missing input.
@@ A.6 Compute resources
All experiments were conducted on NVIDIA RTX A6000 GPUs. Training an MLP model for MNIST and Fashion-MNIST experiments across tasks took less than one minute. Training unsupervised learning models on CIFAR-10 and CIFAR-100 of Section 4.2 took approximately 15 minutes. Training the combined supervised and unsupervised models of Section 4.3 took approximately one hour and 15 minutes. The majority of the compute was spent on hyperparameter tuning. The total training time for hyperparameter tuning of the models of Section 4.1 is ±$\pm 50$h. The total training time for hyperparameter tuning of the models of Section 4.2 is ±$\pm 30$h for MNIST and Fashion-MNIST and ±$\pm 400$h for CIFAR-10 and CIFAR-100. The total training time for hyperparameter tuning of the models of Section 4.3 is ±$\pm 25$h for MNIST and Fashion-MNIST and ±$\pm 750$h for CIFAR-10. The total training time for the tuning of the bimodal models of Section 4.5 is ±$\pm 20$h.
@@ Appendix B Results
Tables 10 to 12 report the result illustrated in Figures 3, 4, 5, and 7. These results are obtained on the test set of the datasets for five different weight initialisations.
Table 10: Classification accuracy and class average image generation for models considered in Section 4.1. Higher accuracy and lower RMSE are better. We report the mean +/- sem over five seeds.
@@
bPC $98.10^{\pm 0.05}$ $89.24^{\pm 0.12}$ $0.0581^{\pm 0.0004}$ $0.0415^{\pm 0.0005}$
hybridPC $86.22^{\pm 0.15}$ $80.34^{\pm 0.11}$ $0.0612^{\pm 0.0003}$ $0.0480^{\pm 0.0016}$
genPC $83.48^{\pm 0.21}$ $77.00^{\pm 0.16}$ $\mathbf{0.0198^{\pm 0.0001}}$ $0.0140^{\pm 0.0001}$
discPC $98.43^{\pm 0.01}$ $\mathbf{89.74^{\pm 0.14}}$ $0.3133^{\pm 0.0224}$ $0.3326^{\pm 0.0024}$
BP $\mathbf{98.48^{\pm 0.10}}$ $89.66^{\pm 0.11}$ $\mathbf{0.0198^{\pm 0.0001}}$ $\mathbf{0.0128^{\pm 0.0001}}$
Table 11: Image reconstruction RMSE from latent representations for models considered in Section 4.2. Lower RMSE is better. We report the mean +/- sem over five seeds.
@@
genPC $0.2473^{\pm 0.0020}$ $0.2868^{\pm 0.0013}$ $0.1837^{\pm 0.0009}$ $0.2077^{\pm 0.0003}$
hybridPC $0.2401^{\pm 0.0012}$ $0.2508^{\pm 0.0007}$ $0.1664^{\pm 0.0015}$ $0.2089^{\pm 0.0071}$
AE $\mathbf{0.1565^{\pm 0.0006}}$ $\mathbf{0.1868^{\pm 0.0001}}$ $0.1135^{\pm 0.0050}$ $0.1171^{\pm 0.0042}$
BP $0.1969^{\pm 0.0004}$ $0.2084^{\pm 0.0002}$ $\mathbf{0.0964^{\pm 0.0004}}$ $\mathbf{0.0983^{\pm 0.0002}}$
Table 12: Classification accuracy and image reconstruction RMSE from latent representations for models considered in Section 4.3. Higher accuracy and lower RMSE are better. We report the mean +/- sem over five seeds.
@@
hybridBP $\mathbf{98.48^{\pm 0.03}}$ $\mathbf{89.61^{\pm 0.13}}$ $\mathbf{90.11^{\pm 0.11}}$ $\mathbf{0.1249^{\pm 0.0004}}$ $\mathbf{0.1707^{\pm 0.0004}}$ $\mathbf{0.1048^{\pm 0.0011}}$
Table 13: Classification accuracy and class average image generation for bimodal genPC and bPC considered in Section 4.5. Higher accuracy and lower RMSE are better. We report the mean +/- sem over five seeds.
@@
bPC $\mathbf{97.80^{\pm 0.05}}$ $\mathbf{89.05^{\pm 0.10}}$ $\mathbf{0.0506^{\pm 0.0007}}$ $\mathbf{0.0431^{\pm 0.0004}}$
genPC $88.28^{\pm 0.10}$ $80.53^{\pm 0.14}$ $0.1561^{\pm 0.0004}$ $0.2027^{\pm 0.0031}$

### v2

@@ 5 Conclusion
We begin with a detailed explanation of the methods underlying the experiments presented in the main paper. Next, we provide summary tables containing the exact values plotted in the figures. We then explore the learning performance of three additional predictive coding models. Following that, we examine the effect of scaling the discriminative and generative energies in bPC on both supervised and unsupervised learning performance. We also analyse the impact of removing max-pooling on generative artefacts. We present samples generated by a combined discBP and genBP model. After, we perform parameter count matched experiments to validate our results. We then discuss the relationship between hybridPC and bPC. Finally, we evaluate whether bPC learns effectively in deep models.
@@ Appendix A Details of experiments
Here, we detail the training setup and the evaluation procedure of the experiments in the paper. The code with all models and experiments will be available upon publication . Our implementation of the predictive coding models was adapted from pcx_.
@@ A.1 Datasets
We evaluate the models on four standard image classification benchmarks: MNIST(lecun2010mnist_), Fashion-MNIST(xiao2017fashionmnist_), CIFAR-10, and CIFAR-100(krizhevsky2009learning_). Below, we summarise their properties and how they were subdivided in our experiments for training, validation and testing. The images of all datasets were normalised so that pixel values varied between 1-1 and $1$.
@@ A.3 Training procedures
Require : Model with neural activities $x$, parameters $\theta$, energy $E$ and initialisation function $init(\cdot)$. Dataset $\{y_{p}\}_{p=1}^{P}$ with $P$ mini-batches of $B$ elements. Number of epochs NN. Activity optimiser $optim_{x}(\cdot)$, number of activity updates $K$, and parameter optimiser $optim_{\theta}(\cdot)$
for $n=1$ to NN do
@@
Require : Model with parameters $\theta$ and forward pass $fp(\cdot,\theta)$. Energy function $E$. Dataset $\{y_{p}\}_{p=1}^{P}$ with $P$ mini-batches of $B$ elements. Number of epochs NN. Parameter optimiser $optim_{\theta}(\cdot)$
for $n=1$ to NN do
@@
The optimisation procedures differ between activity updates and parameter updates. In Algorithm 1, neural activities are updated using stochastic gradient descent with momentum. This optimiser outperforms Adam in PC models (pcx_). In contrast, Algorithm 2 does not include activity updates and only optimises parameters. For parameter updates in both algorithms, we use the AdamW optimiser (loshchilov2017decoupled_). The optimiser hyperparameters were determined via a hyperparameter search.
@@
The hybridBP model includes a stop-gradient operation in its energy function. This ensures that its second loss term is only applied during parameter updates, not iterative inference. This loss term objective is to improve the discriminative initialisation of the model’s latent layer. In the broader machine learning literature, hybridBP can be viewed as a generative model that performs Bayesian inference by combining amortised inference with iterative refinement of the latent variables (marino2018iterative_).
@@ A.4 Training hyperparameters
In this Section, we report the hyperparameter search space for the experiments in Sections 4.1, 4.2, 4.3, and 4.5, as well as the training parameters for the models in Section 4.4. All hyperparameter tuning was performed using Bayesian optimisation from Weights and Biases (wandb_).
Tables 5, 6, and 7 list the hyperparameter search spaces for the models trained in Sections 4.1, 4.2, and 4.3, respectively. The hyperparameter search was conducted using the validation sets of each dataset. In Section 4.1, discPC and genPC were tuned for classification accuracy and generation RMSE separately, while bPC and hybridPC were jointly tuned on both metrics. This ensures a fair comparison, as bPC and hybridPC have twice as many parameters as discPC and genPC. For joint tuning, we combined the two metrics using the objective $2\cdot(1-\text{accuracy}/100)+\text{RMSE}$. In Section 4.2, the models were tuned for the image reconstruction MSE, linear decoding accuracy and generation FID separately. For the reconstruction, the RMSE was reported instead of the MSE for consistency with the generation RMSE. In Section 4.3, all models were jointly optimised for classification accuracy and reconstruction MSE. We combined the two metrics using $2\cdot(1-\text{accuracy}/100)+\text{MSE}$ for MNIST and Fashion-MNIST, and $(1-\text{accuracy}/100)+4\cdot\text{MSE}$ for CIFAR-10. The RMSE was also reported after tuning instead of the MSE. The weightings in the combined objectives compensate for the scale difference between the metrics.
@@ A.5 Evaluation
Classification accuracy We evaluate classification performance using accuracy, defined as the percentage of input images correctly classified by each model.
@@
Conditional generation of mean class images We evaluate the generative performance of the models by computing the root mean squared error (RMSE) between images generated for each class and the corresponding class-average image. We obtain the class-average image by computing the average image across all images belonging to that class in the evaluation set. To compute the RMSE, we calculate the squared error for each pixel between the generated image and the average image for the corresponding class. These errors are averaged across all pixels and classes, and the square root of this mean is reported as the final RMSE.
@@
Image reconstruction from representations We assess the quality of learned representations by measuring the RMSE between reconstructed images and their original inputs. RMSE is computed in the same manner as for conditional image generation.
@@
Linear readout/decoding accuracy We further evaluate the quality of the learned representations by measuring their linear readout accuracy. The representations are obtained in the same way as for image reconstruction. A linear classifier head is then trained with backpropagation for each model to classify images based on their representations. This evaluation tests whether representations of different classes are linearly separable. Higher decoding accuracy reflects better representations.
FID of generated image samples. We compare the ability of bPC, genPC, and hybridPC models to learn probability distributions using stochastic extensions of predictive coding. Predictive coding can be equipped with stochastic neural dynamics to model distributions of sensory inputs in an unsupervised manner by injecting noise into the inference process. This idea was originally proposed in (mcpc) for the genPC model, and here we apply it to genPC, hybridPC, and bPC.
When stochastic dynamics are applied, the noisy inference process can be shown to generate samples from the posterior distribution of the probabilistic model defined by predictive coding, conditioned on the sensory input. These posterior samples can then be used for parameter learning through Monte Carlo Expectation Maximisation.
The training procedure mirrors that of the deterministic experiments: each iteration consists of several (now noisy) activity updates followed by a single parameter update. This is equivalent to using a single posterior sample per parameter update, a common practice in related generative models such as variational autoencoders.
To accelerate inference, we further incorporate momentum into the neural dynamics, as introduced for predictive coding in (pcx). The resulting dynamics correspond to a discrete-time second-order Langevin process. In practice, this is implemented by adding Gaussian noise to the gradients of the energy function and updating activities with stochastic gradient descent with momentum:
$\Delta x_{l}=lr_{x}r_{l},$ (7)
$\Delta r_{l}=-lr_{x}\nabla_{x_{l}}E;-;lr_{x}(1-m)r_{l};+;\sqrt{2(1-m)lr_{x}},N.$ (8)
After training, the models can be used to generate input samples. For a top-down predictive coding model, this is done by ancestral sampling: first sampling the top latent layer $x_{L}\sim N(0,I)$, and then recursively sampling each lower layer from its conditional Gaussian distribution $N(x_{l};f_{l}(x_{l+1},W_{l+1}),I)$. The prior distribution on $x_{L}$ emerges from the additional activity decay applied to $x_{L}$ for the unsupervised learning experiments.
To evaluate generative performance, we compute the Fréchet Inception Distance between generated and real samples. We use the open-source library pytorch-fid (Seitzer2020FID), adapted to MNIST by replacing the standard Inception network with a ResNet-18 trained on MNIST as the feature extractor.
As a baseline, we include a variational autoencoder (VAE) trained on the same learning task. The VAE architecture and parameter count are matched to those of hybridPC and bPC to enable a fair comparison. Training follows the standard VAE procedure: the encoder network approximates the posterior distribution over latent variables, while the decoder parametrises the generative model. The model is optimised end-to-end via backpropagation using the standard VAE objective.
@@
Visualisation of Energy Landscapes on XOR We visualise the energy landscapes of trained genPC, discPC, and bPC models on the XOR task. To generate the landscape, we clamp the input layer neurons to a 2D coordinate within the range [-3,3], sampled at intervals of 0.25 along both axes. Simultaneously, we clamp the output layer $x_{L}$ to the one-hot encoding of one of the two classes. For each coordinate–label combination, we run 10,000 steps of iterative inference to ensure convergence to equilibrium and record the final energy of the model. This procedure is repeated over the grid of 2D inputs and both class labels, allowing us to plot the full energy landscape. For fair comparison, we also include the combined energy landscape of discPC and genPC. Together, the combined parameter count of discPC and genPC matches the total parameter count of bPC. Equal weighting is applied to the energy values of the discPC and genPC models when adding their energy, as we use equal weighting for the discriminative and generative components in the bPC model ($\alpha_{disc}=\alpha_{gen}=1$).
@@
MNIST image generation with PC models We evaluate the generative capabilities of the bPC model and a combined genPC and discPC model on the MNIST dataset. The goal is to assess how likely these models are to assign low energy to implausible, label-inconsistent samples.
@@
  • Fréchet Inception Distance (FID): The FID measures the distance between the distributions of real and generated images (HeuselRUNKH17_). We use the public implementation of FID from Seitzer2020FID_, but modify it to use a ResNet classifier trained on MNIST, instead of the original Inception model trained on ImageNet. This ensures that the FID score better reflects visual quality and diversity on the MNIST domain.
  • Inception Score (IS): The Inception Score evaluates how easily a classifier can identify the class of a generated image (salimans2016improvedtechniquestraininggans_). We used a publicly available implementation for the MNIST dataset (inception_).
@@
To assess model robustness to missing input data, we evaluate classification accuracy under varying levels of input occlusion. This evaluation follows the same procedure as standard classification (described above), but for images missing a random subset of pixels. The proportion of missing pixels ranges from 10% to 90%, in increments of 10%. Missing pixels are selected uniformly at random, independently of their spatial location within the image. For all models, missing pixels in $x_{1}$ are initialised to zero activity and are left unclamped, allowing the model to update their values during inference. We repeat the same classification procedure as before, but we increased the number of inference steps to 600,000 to ensure convergence. This is necessary because predictive coding models converge more slowly when input information is incomplete (tang2023recurrent_). In addition to measuring classification accuracy, we also record the post-inference activity of neurons in $x_{1}$ to qualitatively assess the model’s ability to fill in the missing input.
@@ A.6 Compute resources
All experiments were conducted on NVIDIA RTX A6000 GPUs. Training an MLP model for MNIST and Fashion-MNIST experiments across tasks took less than one minute. Training unsupervised learning models on CIFAR-10 and CIFAR-100 of Section 4.2 took approximately 15 minutes. Training the combined supervised and unsupervised models of Section 4.3 took approximately one hour and 15 minutes. The majority of the compute was spent on hyperparameter tuning. The total training time for hyperparameter tuning of the models of Section 4.1 is $\pm 50$h. The total training time for hyperparameter tuning of the models of Section 4.2 is $\pm 30$h for MNIST and Fashion-MNIST and $\pm 400$h for CIFAR-10 and CIFAR-100. The total training time for hyperparameter tuning of the models of Section 4.3 is $\pm 25$h for MNIST and Fashion-MNIST and $\pm 750$h for CIFAR-10. The total training time for the tuning of the bimodal models of Section 4.5 is $\pm 20$h.
A.7 LLM use
ChatGPT Edu was used to polish writing.
@@ Appendix B Results
Tables 10 to 15 report the result illustrated in Figures 3, 4, 5, and 7. These results are obtained on the test set of the datasets for five different weight initialisations.
Table 10: Classification accuracy and class average image generation for models considered in Section 4.1. Higher accuracy and lower RMSE are better. We report the mean +/- sem over five seeds. Results indicated with are significantly worse in performance than bPC determined using an independent-samples t-test (n=5, $p<0.05$).
@@
bPC $98.10^{\pm 0.05}$ $89.24^{\pm 0.12}$ $0581^{\pm 0.0004} 0.0$$0415^{\pm 0.0005}$
hybridPC $86.22^{\pm 0.15*}$ $80.34^{\pm 0.11*}$ $0612^{\pm 0.0003} 0.0$$0480^{\pm 0.0016}$
genPC $83.48^{\pm 0.21*}$ $77.00^{\pm 0.16*}$ $0.0198^{\pm 0.0001}$} $0.0140^{\pm 0.0001}$
discPC $98.43^{\pm 0.01}$ $89.74^{\pm 0.14}$} $0.3133^{\pm 0.0224*}$ $0.3326^{\pm 0.0024*}$
BP $98.48^{\pm 0.10}$} $89.66^{\pm 0.11}$ $0.0198^{\pm 0.0001}$} $0.0128^{\pm 0.0001}$}
Table 11: Image reconstruction RMSE from latent representations for models considered in Section 4.2. Lower RMSE is better. We report the mean +/- sem over five seeds. Results indicated with are significantly worse in performance than bPC determined using an independent-samples t-test (n=5, $p<0.05$).
@@
genPC $0.2473^{\pm 0.0020*}$ $0.2868^{\pm 0.0013*}$ $0.1837^{\pm 0.0009*}$ $0.2077^{\pm 0.0003*}$
hybridPC $0.2401^{\pm 0.0012*}$ $0.2508^{\pm 0.0007}$ $0.1664^{\pm 0.0015*}$ $0.2089^{\pm 0.0071*}$
AE $0.1565^{\pm 0.0006}$} $0.1868^{\pm 0.0001}$} $0.1135^{\pm 0.0050}$ $0.1171^{\pm 0.0042}$
BP $0.1969^{\pm 0.0004}$ $2084^{\pm 0.0002} 0.0$$0964^{\pm 0.0004}$} $0.0983^{\pm 0.0002}$}
Table 12: Linear decoding accuracy (%) across datasets for different models considered in Section 4.2. Higher is better. We report the mean $\pm$ sem over five seeds. Results indicated with are significantly worse in performance than bPC determined using an independent-samples t-test (n=5, $p<0.05$).
Model MNIST Fashion-MNIST CIFAR-10 CIFAR-100
bPC $89.99^{\pm 0.01}$ $81.82^{\pm 0.01}$ $50.52^{\pm 0.01}$ $60.68^{\pm 0.01}$
genPC $86.93^{\pm 0.01*}$ $80.44^{\pm 0.01*}$ $48.23^{\pm 0.02*}$ $50.45^{\pm 0.02*}$
hybridPC $89.07^{\pm 0.01*}$ $81.59^{\pm 0.01}$ $49.61^{\pm 0.01*}$ $59.31^{\pm 0.01*}$
AE $90.76^{\pm 0.01}$ $83.31^{\pm 0.01}$ $47.29^{\pm 0.02}$ $63.78^{\pm 0.04}$
hybridBP $93.44^{\pm 0.01}$} $78.69^{\pm 0.018*}$ $50.63^{\pm 0.01}$} $85.14^{\pm 0.01}$}
Table 13: FID scores for different models trained with 50 and 250 activity updates before each weight update considered in Section 4.2. Lower is better. We report the mean $\pm$ sem over five seeds. The VAE only has one value because it does not have iterative inference.
Model FID @ 50 updates FID @ 250 updates
bPC $5.21^{\pm 0.26}$ $3.34^{\pm 0.53}$
genPC $7.86^{\pm 1.27}$ $4.56^{\pm 0.30}$
hybridPC $5.01^{\pm 0.30}$ $4.28^{\pm 0.32}$
VAE $5.79^{\pm 0.21}$
Table 14: Classification accuracy and image reconstruction RMSE from latent representations for models considered in Section 4.3. Higher accuracy and lower RMSE are better. We report the mean +/- sem over five seeds. Results indicated with are significantly worse in performance than bPC determined using an independent-samples t-test (n=5, $p<0.05$).
@@
hybridBP $98.48^{\pm 0.03}$} $89.61^{\pm 0.13}$} $90.11^{\pm 0.11}$} $0.1249^{\pm 0.0004}$} $0.1707^{\pm 0.0004}$} $0.1048^{\pm 0.0011}$}
Table 15: Classification accuracy and class average image generation for bimodal genPC and bPC considered in Section 4.5. Higher accuracy and lower RMSE are better. We report the mean +/- sem over five seeds. Results indicated with are significantly worse in performance than bPC determined using an independent-samples t-test (n=5, $p<0.05$).
@@
bPC $97.80^{\pm 0.05}$} $89.05^{\pm 0.10}$} $0.0506^{\pm 0.0007}$} $0.0431^{\pm 0.0004}$}
genPC $88.28^{\pm 0.10*}$ $80.53^{\pm 0.14*}$ $0.1561^{\pm 0.0004*}$ $0.2027^{\pm 0.0031*}$